OpenStudy (anonymous):

If (x, 1/100) lies on the graph of y = 10^x, then x =

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

substitute the given point into the equation

OpenStudy (anonymous):

(x, 1/100) lies on the graph so corresponding y to this x is 1/100\[\frac{1}{100}=10^x\]

OpenStudy (anonymous):

hes lieing to you, the correct answer is $$x=-2$$

OpenStudy (anonymous):

don't trust tushara